Hi can you help me solve this equation by substitution
3b+2a=2
-2b+a=8

Answer by chibisan(131)


3b+2a=2 (eqn 1)
-2b+a=8 (eqn 2)
from eqn 2;
a=8+2b
Subst. eqn2 to eqn1

3b+2(8+2b)=2
3b+16+4b=2
7b=2-16
7b=-14
b=-2

since b=-2 ;
a=8+(-2)
a=6